Question 27 Determine the Center of the Ellipse (X-4)2 + (Y+3)2 - 1 49 64 A. (4,3) B. (4,-3) C. (7,8) D. (-4,3) A B C D
Answer #1

Option B (4,-3) is correct.

For the ellipse with the equation,

\frac{(x-p)^2}{a^2}+\frac{(y-q)^2}{b^2}=1, the center is always located at the coordinate (p,q).

So looking at hte equation \frac{(x-4)^2}{49}+\frac{(y+3)^2}{64}=1 , we can say that the centre is at (4,-3).